OpenStudy (anonymous):

I=4/3Ma solve for 'a' showing all workings

OpenStudy (anonymous):

\[I = \frac{ 4 }{ 3 } M*a\] \[I \div \frac{ 4 }{ 3 } = M*a\] \[I * \frac{ 3 }{ 4 } = M*a\] \[(I * \frac{ 3 }{ 4 }) \div M = a\]

OpenStudy (anonymous):

Thx